# DOMAIN: milDly cOnservative Model-bAsed offlINe reinforcement learning

### v2

@@ V-D The Parameter Analysis
This part gives further analysis for weight $\lambda$ and rollout length $H$. Table VIII analyzes the impact of two parameter changes on the performance of the agent. Due to high training cost, here we fix one parameter and test the impact of changes in another parameter on agent performance.
The effect of weight $\lambda$. The parameter $\lambda$ is used to regularization between Bellman error and $Q$ value of data, and guarantee the lower bound of $Q$-function for inaccuracy model data. A larger $\lambda$ is needed to guarantee the lower bound of inaccuracy model data in narrower data distribution. However, for diverse data distribution, too large $\lambda$ may cause the lower bound to be too loose, thus reducing agent performance. Table VIII shows that in narrower “Medium” and “Medium-Expert” datasets, too small $\lambda$ leads to unstable training, thus reducing the agent performance. For diverse “Medium-Replay” datasets, too large $\lambda$ causes the high constraint, thus limiting the improvement of the agent.
The effect of rollout length $H$. When rollout length $H=1$, the trained environment model has high prediction accuracy. As $H$ increases, the accuracy of prediction decreases, but it is beneficial to explore the OOD region. If the model data error is too large, the algorithm will not be able to achieve a balance between accurate offline data and inaccurate model data, leading to unstable agent training. Therefore, it is necessary to ensure that the model data error is not too large. Table VIII shows that, as $H$ increases, the agent performance shows a downward trend under the Walker2d task since the error of the trained environment is large. Under the Halfcheetah and Hopper tasks, as $H$ increases, the agent performance does not show a downward trend, but the optimal $H$ is different for different datasets.
@@ VI Conclusion
This paper proposes a mildly conservative model-based offline RL (DOMAIN) algorithm in this paper. By designing a model data adaptive sampling distribution, this algorithm can dynamically adjust the model data penalties, increasing the penalty for model data with larger errors while reducing the penalty for accurate model data, thereby reducing its conservatism. Moreover, DOMAIN avoids unreliable estimation of model uncertainty. Theoretical analysis demonstrates that DOMAIN can guarantee safe policy improvements and $Q$-values learned by DOMAIN outside the region are lower bounds of the true $Q$-values. Experimental results on the D4RL benchmark show that DOMAIN outperforms previous RL algorithms on most datasets and achieves the best generalization performance in tasks that require adaptation to unseen behaviors.
Future works can focus on the below aspects: 1) Since incorporating conservatism during the training process constrains the improvement of agent performance, future efforts can focus on optimizing the environment model itself to further enhance the overall performance. 2) Since training the discriminator to judge the model data error has a certain error, future works can focus on finding a more accurate and stable way to measure the model data error. 3) Most RL algorithms need to tune parameters constantly for different datasets. The less parameter tuning RL methods need further studied in the future.
@@ A-A Proof of Theorem 1
$+f\left\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T_{\mathcal{M}}}[\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})]\right\}$
@@
$\|\hat{\mathcal{T}}^{\pi}Q_{1}-\hat{\mathcal{T}}^{\pi}Q_{2}\|_{\infty}$ (19)
@@ A-B Proof of Theorem 2
The difference between empirical MDP $\overline{\mathcal{M}}$ and actual MDP ${\mathcal{M}}$ lies in the sampling error. The sampling error comes from the inadequate dataset (it is difficult for the dataset to cover the entire data space). Therefore, if the dataset is large enough, the sampling error would be smaller and the gap between empirical MDP and actual MDP would become smaller. Therefore, we make the following assumptions following prior works [13, 21, 31, 34] :
@@
$\left|\left(\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}\right)-\left(\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right)\right|<\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\sqrt{|\mathcal{D}(s,a)|}}$ (21)
@@
$=$ $\mathcal{T}^{\pi}\widehat{Q}^{k}+\underbrace{(1-f)\Delta_{1}(s,a)+{\color[rgb]{0,0,0}f\Delta_{2}(s,a)-\lambda\eta(s,a)}}_{:=\Delta(s,a)}$
@@
Note that the error of data from OOD region is likely to be greater than that from offline region, that is, $\omega(s,a)$ become larger and $d(s,a)$ become smaller for model data in OOD region, thereby satisfying $\Delta_{3}>0$. When the offline data is sufficient enough, the $\Delta_{1}$ and $\Delta_{2}$ approach zero. Therefore, the term $\Delta<0$ holds with high probability, thus $\left(I-\gamma P^{\pi}\right)^{-1}\mathbb{E}_{a\sim\pi(a\mid s)}\left[\Delta(s,a)\right]<0$ holds with high probability for OOD data.
@@
$=$ ${\color[rgb]{0,0,0}\int_{\mathcal{A}}\frac{\omega(s,a)-d(s,a)}{(1-f)d^{\pi_{b}}(s)\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}+fd_{\mathcal{\widehat{\mathcal{M}}}}^{\pi}(s)}\mathrm{d}a}$
$\geq$ $\frac{{\color[rgb]{0,0,0}\int_{\mathcal{A}}\omega(s,a)\mathrm{d}a}}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}$
$-\frac{{\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{\mathcal{M}}}}^{\pi}(s)}$
@@
${\color[rgb]{0,0,0}\int_{\mathcal{A}}\omega(s,a)\mathrm{d}a}$ (30)
$>$ $\frac{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}{\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}$
$=$ $\xi(s){\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}$
Substituting Eq. (30) into (28), it can further derive:
$\lambda\cdot\min_{s}\frac{(\xi(s)-1){\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}$ (31)
@@
$\Rightarrow$ $\lambda\geq\frac{(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\}}}{\min_{s}\left\{\frac{(\xi(s)-1){\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
$+\frac{f[\max\limits_{s,a}\left\{\left|r_{\mathcal{M}}-r_{\mathcal{M}}\right|\right\}+\frac{2\gamma R_{\max}}{1-\gamma}\max\limits_{s,a}\left\{D_{TV}\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\right\}]}{\min_{s}\left\{\frac{(\xi(s)-1){\color[rgb]{0,0,0}\int_{\mathcal{A}}d(s,a)\mathrm{d}a}}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
@@ A-C Proof of Theorem 3
$=$ $\mathbb{E}_{s\sim d(s)}\left[{\color[rgb]{0,0,0}\int_{\mathcal{A}}\frac{\pi(a\mid s)\left[\rho(s,a)-\omega(s,a)\right]}{(1-f)d^{\pi_{b}}(s)\pi_{b}(a\mid s)+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a\mid s)}\mathrm{d}a}\right]$
$\geq$ $\mathbb{E}_{s\sim d(s)}\left[\frac{{\color[rgb]{0,0,0}\int_{\mathcal{A}}(\rho(s,a)-\omega(s,a))\mathrm{d}a}}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s})\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}\right]$
@@ A-D Proof of Theorem 4
$=$ $\frac{1}{1-\gamma}\left|{\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{A}}\Big{[}d_{\overline{\mathcal{M}}}^{\pi}(s)\pi(a|s)r_{\overline{\mathcal{M}}}-d_{\mathcal{M}}^{\pi}(s)\pi(a|s)r_{\mathcal{M}}\Big{]}\mathrm{d}s\mathrm{d}a}\right|$
$\leq$ $\frac{1}{1-\gamma}\left|{\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{A}}d_{\overline{\mathcal{M}}}^{\pi}(s)\pi(a|s)\big{[}r_{\overline{\mathcal{M}}}(s,a)-r_{\mathcal{M}}(s,a)\big{]}\mathrm{d}s\mathrm{d}a}\right|$
$+\frac{1}{1-\gamma}\left|{\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{A}}\left(d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}(s,a)\mathrm{d}s\mathrm{d}a}\right|$
@@
$\left|{\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{A}}d_{\overline{\mathcal{M}}}^{\pi}\pi\left[r_{\overline{\mathcal{M}}}-r_{\mathcal{M}}\right]\mathrm{d}s\mathrm{d}a}\right|\leq\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}}\left[\frac{C_{r,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (35)
@@
$\leq$ ${\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{S^{\prime}}}\left|\int_{\mathcal{A}}\left(T_{\overline{\mathcal{M}}}-T_{\mathcal{M}}\right)\pi\mathrm{d}a\right|d_{\overline{\mathcal{M}}}^{\pi}(s)\mathrm{d}s\mathrm{d}s^{\prime}}$
@@
$\left|{\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{A}}\left[d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right]\pi(a|s)r_{\mathcal{M}}(s,a)\mathrm{d}s\mathrm{d}a}\right|$ (37)
@@
$\leq$ $\frac{1}{1-\gamma}\left|{\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{A}}d_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a\mid s)\left[r_{\widehat{\mathcal{M}}}(s,a)-r_{\mathcal{M}}\right]\mathrm{d}s\mathrm{d}a}\right|$
$+\frac{1}{1-\gamma}\left|{\color[rgb]{0,0,0}\int_{\mathcal{S}\times\mathcal{A}}\left(d_{\widehat{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}\mathrm{d}s\mathrm{d}a}\right|$

### v3

@@ V-D The Parameter Analysis
This part gives further analysis for weight $\lambda$ and rollout length $H$. Table V analyzes the impact of two parameter changes on the performance of the agent. Due to high training cost, we fix one parameter and test the impact of changes in another parameter on agent performance.
The effect of weight $\lambda$. The parameter $\lambda$ is used to regularization between Bellman error and $Q$ value of data, and guarantees the lower bound of $Q$-function for inaccuracy model data. A larger $\lambda$ is needed to guarantee the lower bound of inaccuracy model data in narrower data distribution. However, for diverse data distribution, too large $\lambda$ may cause the lower bound to be too loose, thus reducing agent performance. Table V shows that in narrower “Medium” and “Medium-Expert” datasets, too small $\lambda$ leads to unstable training, thus reducing the agent performance. For diverse “Medium-Replay” datasets, too large $\lambda$ causes the high constraint, thus limiting the improvement of the agent.
The effect of rollout length $H$. When rollout length $H=1$, the trained environment model has high prediction accuracy. As $H$ increases, the accuracy of prediction decreases, but it is beneficial to explore the OOD region. If the model data error is too large, the algorithm will not be able to achieve a balance between accurate offline data and inaccurate model data, leading to unstable agent training. Therefore, it is necessary to ensure that the model data error is not too large. Table V shows that, as $H$ increases, the agent performance shows a downward trend under the Walker2d task since the error of the trained environment is large (Fig. 3). For Halfcheetah and Hopper tasks, the environment error is small. Therefore, as $H$ increases, the agent performance does not show a downward trend, but the optimal $H$ is different for different datasets.
@@ VI Conclusion
This paper proposes a mildly conservative model-based offline RL (DOMAIN) algorithm. By designing a model data adaptive sampling distribution, DOMAIN can dynamically adjust the model data penalty, increasing the penalty for inaccuracy model data while reducing the penalty for accurate model data, thereby reducing its conservatism. Moreover, DOMAIN avoids unreliable estimation of model uncertainty. Theoretical analysis demonstrates that DOMAIN can guarantee safe policy improvements and $Q$-values learned by DOMAIN outside the region are lower bounds of the true $Q$-values. Experimental results on the D4RL benchmark show that DOMAIN outperforms previous RL algorithms on most datasets. Since incorporating conservatism during the training process constrains the improvement of agent performance, future efforts can focus on optimizing the environment model itself to further enhance the overall performance.
@@ A-A Proof of Theorem 1
$+f\left\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}[\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})]\right\}$
@@
$\|\widehat{\mathcal{T}}^{\pi}Q_{1}-\widehat{\mathcal{T}}^{\pi}Q_{2}\|_{\infty}$ (19)
@@ A-B Proof of Theorem 2
Before proving Theorem 2, we make the following assumptions following prior works [13, 21, 31, 34]:
@@
$\left|\left(\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}\right)-\left(\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right)\right|<\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\sqrt{|\mathcal{D}(s,a)|}}$
@@
$=$ $\mathcal{T}^{\pi}\widehat{Q}^{k}+\underbrace{(1-f)\Delta_{1}(s,a)+f\Delta_{2}(s,a)-\lambda\eta(s,a)}_{:=\Delta(s,a)}$
@@
$=$ $\sum_{a}\frac{\omega(s,a)-d(s,a)}{(1-f)d^{\pi_{b}}(s)\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}+fd_{\mathcal{\widehat{\mathcal{M}}}}^{\pi}(s)}$
$\geq$ $\frac{\sum_{a}\omega(s,a)}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}$
$-\frac{\sum_{a}d(s,a)}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{\mathcal{M}}}}^{\pi}(s)}$
@@
$\sum_{a}\omega(s,a)$
$>$ $\frac{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}\sum_{a}d(s,a)$
$=$ $\xi(s)\sum_{a}d(s,a)$
Substituting Eq. (A-B) into (27), it can further derive:
$\lambda\cdot\min_{s}\frac{(\xi(s)-1)\sum_{a}d(s,a)}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}$
@@
$\Rightarrow$ $\lambda\geq\frac{(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\}}}{\min_{s}\left\{\frac{(\xi(s)-1)\sum_{a}d(s,a)}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
$+\frac{f[\max\limits_{s,a}\left\{\left|r_{\mathcal{M}}-r_{\mathcal{M}}\right|\right\}+\frac{2\gamma R_{\max}}{1-\gamma}\max\limits_{s,a}\left\{D_{TV}\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\right\}]}{\min_{s}\left\{\frac{(\xi(s)-1)\sum_{a}d(s,a)}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
@@ A-C Proof of Theorem 3
$=$ $\mathbb{E}_{s\sim d(s)}\left[\sum_{a}\frac{\pi(a\mid s)\left[\rho(s,a)-\omega(s,a)\right]}{(1-f)d^{\pi_{b}}(s)\pi_{b}(a\mid s)+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a\mid s)}\right]$
$\geq$ $\mathbb{E}_{s\sim d(s)}\left[\frac{\sum_{a}(\rho(s,a)-\omega(s,a))}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s})\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}\right]$
@@ A-D Proof of Theorem 4
$=$ $\frac{1}{1-\gamma}\left|\sum_{s,a}\Big{[}d_{\overline{\mathcal{M}}}^{\pi}(s)\pi(a|s)r_{\overline{\mathcal{M}}}-d_{\mathcal{M}}^{\pi}(s)\pi(a|s)r_{\mathcal{M}}\Big{]}\right|$
$\leq$ $\frac{1}{1-\gamma}\left|\sum_{s,a}d_{\overline{\mathcal{M}}}^{\pi}(s)\pi(a|s)\big{[}r_{\overline{\mathcal{M}}}(s,a)-r_{\mathcal{M}}(s,a)\big{]}\right|$
$+\frac{1}{1-\gamma}\left|\sum_{s,a}\left(d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}(s,a)\right|$
@@
$\left|\sum_{s,a}d_{\overline{\mathcal{M}}}^{\pi}\pi\left[r_{\overline{\mathcal{M}}}-r_{\mathcal{M}}\right]\right|\leq\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}}\left[\frac{C_{r,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (31)
@@
$\leq$ $\sum_{s,s^{\prime}}\left|\int_{\mathcal{A}}\left(T_{\overline{\mathcal{M}}}-T_{\mathcal{M}}\right)\pi\right|d_{\overline{\mathcal{M}}}^{\pi}(s)$
@@
$\left|\sum_{s,a}\left[d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right]\pi(a|s)r_{\mathcal{M}}(s,a)\right|$ (33)
@@
$\leq$ $\frac{1}{1-\gamma}\left|\sum_{s,a}d_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a\mid s)\left[r_{\widehat{\mathcal{M}}}(s,a)-r_{\mathcal{M}}\right]\right|$
$+\frac{1}{1-\gamma}\left|\sum_{s,a}\left(d_{\widehat{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}\right|$
@@
$\left|J\left(\mathcal{M}_{f},\pi\right)-J(\mathcal{M},\pi)\right|$